OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1- 8 - 3x = 5x + 16 Equation #2- 9x - 14 = 8x - 15 A. The solution to equation #1 is larger than the solution to equation #2. B. The solution to equation #1 is smaller than the solution to equation #2. C. The solution to equation #1 is the same as the solution to equation #2. D. None of the statements above describe solutions to equations shown. @johnweldon1993

OpenStudy (johnweldon1993):

Is that a -8 and a -9x right in front in the 2 equations?

OpenStudy (anonymous):

no not a negative its just a lne that shows you its that equation

OpenStudy (johnweldon1993):

Ahh thank you...hang on ...

OpenStudy (anonymous):

Ok no problem

OpenStudy (anonymous):

The set of ordered pairs below defines a relation. In each ordered pair (x, y), x represents the daily wages of an electrician, and y represents the daily wages of her assistant. {(95, 64), (110, 42), (80, 42), (125, 80), (110, 80)} What is the difference between the values of the least and greatest elements in the domain of the relation? 15 30 38 45

OpenStudy (anonymous):

Which equation is not equivalent to the formula m = ca? c equals m over a a equals c over m a equals m over c m = ac

OpenStudy (anonymous):

Simplify quantity 2 times x plus 10 all over 2 2x + 5 x + 10 x + 5 x

OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)? 1 over 5 5 negative 1 over 5 -5

OpenStudy (anonymous):

ake recorded the number of heartbeats of rabbits for different time intervals, on a scatter plot. He drew a line of best fit on the scatter plot, as shown below.

OpenStudy (anonymous):

Which statement best describes the meaning of the slope as a rate of change for this line of best fit? It represents the increase in the heartbeat rate of rabbits. It represents the change in heartbeat rate for unit change in time. It represents the average heartbeat rate of rabbits, in beats per minute. It represents the time in which a rabbit completes a heartbeat cycle

OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12? A. slope = negative 3 over 2; y-intercept = (0, 6) B. slope = -3; y-intercept = (0, 12) C. slope = negative 2 over 3; y-intercept = (0, 6) D. slope = 6; y-intercept = (0, negative 3 over 2)

OpenStudy (anonymous):

Simplify 2(10u - 4). A. 20u - 4 B. 12u - 6 C. 20u + 8 D. 20u - 8

OpenStudy (anonymous):

Choose the graph below that correctly represents the equation 4x + 3y = 12.

OpenStudy (anonymous):

Connie needs to burn at least 250 calories at the gym today. She can burn 22 calories per minute on the treadmill and 10 calories per minute on the elliptical trainer. Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 250 calories? A. 22t - 10e greater than or equal to 250 B. 22t + 10e less than or greater to 250 C. 22t + 10e greater than or equal to 250 D, 22t - 10e less than or greater to 250

OpenStudy (anonymous):

the ones that say greater to is suppose to be equal too sorry bout that

OpenStudy (anonymous):

Solve for y: 2 over 3y + 2 = 8 A, y = 11 B. y = -11 C. y = -9 D. y = 9

OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? A. y = 3x + 5 B. y = 3x + 3 C. y = 3x - 17 D. y = 3x - 5

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x < -12 or x + 3 greater than or equal to 12 A. A number line with an open circle on -6, shading to the left, and a closed circle on 9, shading to the right. B. A number line with a closed circle on -6, shading to the left, and an open circle on 9, shading to the right. C. A number line with an open circle on -6, a closed circle on 9, and shading in between. D. A number line with a closed circle on -6, an open circle on 9, and shading in between.

OpenStudy (anonymous):

Davenport Music Academy held workshops on classical music composers Beethoven, Strauss, and Vivaldi. The Venn diagram shows the number of seniors who chose to attend workshops on each of the three composers.

OpenStudy (anonymous):

If the total number of seniors in the academy is 55, what is the total number of students who belong to the set (B ∩ ~V) ∩ ~S? 27 19 11 8

OpenStudy (anonymous):

Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + one half (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? 30 ml 35 ml 40 ml 45 ml

OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality 2x + y > 4. A. Dashed line with shading on the side that includes the origin. B. Solid line with shading on the side that does not include the origin. C. Dashed line with shading on the side that does not include the origin. D. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

Choose the graph that correctly corresponds to the equation x = -4.

OpenStudy (anonymous):

Which relation does NOT represent a function? A. (-4, 2), (1, 5), (-2, 0), (1, 6) B. (-5, 3), (-3, 1), (0, -4), (3, -6) C.(2, 0), (4, 6), (6, 8), (10, 6) D. (-3, 6), (-1, 1), (0, 1), (3, 7)

OpenStudy (anonymous):

If C = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12} and D = {numbers divisible by 2 from 1 to 10}, what is CintersectionD? A.{2, 4, 6, 8, 10} B. {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12} C. {1, 3, 5, 7, 9, 11} D.{2, 4, 6, 8, 10, 12}

OpenStudy (anonymous):

Simplify the following expression. 5x - 6y + 5z - x - 12y + 2z A. 6x - 18y + 7z B. 4x - 18y + 7z C. 4x - 6y + 7z D. 6x - 6y + 7z

OpenStudy (anonymous):

Solve for x: quantity x minus 7 all over 3 equals 10 20 37 23 16

OpenStudy (anonymous):

Solve for x: -3(x - 3) = 21 x = -10 x = 10 x = -4 x = 4

OpenStudy (anonymous):

Which statement is correct about the Venn Diagram shown below?

OpenStudy (anonymous):

Everyone surveyed liked cereal or pancakes. A total of 20 people were surveyed. 16 people surveyed liked cereal. 23 people surveyed liked cereal and pancakes.

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? A. x - y = -1 B. x - y = -2 C. x - y = - 5 D. x - y = - 6

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 7 > -14 and 5x less than or greater to 25 A. A number line with an open circle on -7, shading to the left, and a closed circle on 5, shading to the right. B. A number line with a closed circle on -7, shading to the left, and an open circle on 5, shading to the right. C. A number line with a closed circle on -7, an open circle on 5, and shading in between. D. A number line with an open circle on -7, a closed circle on 5, and shading in between.

OpenStudy (anonymous):

Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 (2 points): Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (anonymous):

What is the x-intercept and y-intercept of the graph of 3x - 4y = 12? A. x-intercept = (-4, 0); y-intercept = (0, 3) B. x-intercept = (4, 0); y-intercept = (0, -3) C. x-intercept = (0, -3); y-intercept = (4, 0) D. x-intercept = (0, 3); y-intercept = (-4, 0)

OpenStudy (anonymous):

Solve for x: x - 2 > 2x + 12 A. x > 14 B. x < 14 C.x > -14 D. x < -14